During a breath, the average human lung expands by about 0.50 L.

If this expansion occurs against an external pressure of 1.0 atm, how much work (in J) is done during the expansion?

Express your answer using two significant figures.
Answer #1

Work done W = -P*DV

= -1*0.5 = 0.5 lit.atm

=- 0.5*101 =- 50.66 joule